2. a) Find the derivatives of the following functions with respect to x.



x ^ 3 + y ^ 3 = 3


y = (sin x) ^ tan x



b) Evaluate the 2 ^ (nd) order partial derivatives partial^ 2 u partial x^ 2 and partial^ 2 u partial y^ 2 if u=2x^ 3 +3x^ 2 y+xy^ +y^ .



1
Expert's answer
2021-12-15T08:49:54-0500

 "\begin{aligned}

&x^{3}+y^{3}=3 \\

&y^{3}=3-x^{3}

\end{aligned}"

Differentiating implicitly, we get

"\begin{aligned}

&3 y^{2} \frac{d y}{d x}=-3 x^{2} \\

&\frac{d y}{d x}=-\frac{x^{2}}{y^{2}} \\

&y=\sin (x)^{\tan (x)}

\end{aligned}"

We take the natural logarithm to obtain "\ln y=\tan x \ln \sin x"

Next, we take the exponential of both sides to obtain

"\begin{aligned}

&\frac{1}{y} \cdot \frac{d y}{d x}=\sec ^{2} x \ln \sin x+\tan x \frac{1}{\sin x} \cdot \cos x \\

&=1+\ln \sin x \sec ^{2} x

\end{aligned}"

  "\Longrightarrow \frac{1}{y} \cdot \frac{d y}{d x}=1+\ln \sin x \sec ^{2} x"

Multiplying both sides by "y=\sin x^{\tan x}" , we have

"\begin{aligned}

&\frac{d\left(\sin x^{\tan x}\right)}{d x}=\left(1+\ln \sin x \sec ^{2} x\right) \sin x^{\tan x} \\

&u=2 x^{3}+3 x^{2} y+x y^{2}+y^{2}\end{aligned}"

The 2nd order partial derivatives is given by "u_{x x}, u_{x y}, u_{y y}"

"\begin{aligned}

&u_{x x}=12 x+6 y \\

&u_{x y}=6 x+2 y \\

&u_{y x}=6 x+2 y \\

&u_{y y}=2 x+2

\end{aligned}"